Light is gathered from a distant galaxy and one of the spectral lines of Calcium is observed at 413.26 nm when it should be 393.37 nm.

(a) Find the velocity of this galaxy (hint: Doppler formula).

(b) Find the distance to this galaxy (hint: Hubble Law).
Answer #1

Solution:

According to Doppler shift: dw/w = v/c,

where dw is the wavelength shift,

   w the nominal wavelength,

   v the speed of the object and

   c the speed of light.

Here, we have: v/c = dw/w = (413.26 - 393.37) / 393.37 = 0.051

Therefore v= 0.051 x 300,000 km/s = 15,170 km/s [Part-(A) Ans.]

Now, According to Hubble's law v=H d,

where H = 71 km/s / Mpc
So the galaxy is at a distance of 15,170 / 71 = 214 Mpc